1. Consider a risky portfolio. The end-of-year cash flow derived from the portfolio will be either $75,000 with a probability of 25%, $125,000 with a probability of 50%, or $200,000 with a probability of 25%. The alternative risk-free investment in T-bills pays 3% per year.

a. What is the Sharpe ratio of the portfolio?

b. If you require a risk premium of 8%, how much will you be willing to pay for the portfolio?
Answer #1

1.
Expected returns=25%*(75000/100000-1)+50%*(125000/100000-1)+25%*(200000/100000-1)=31.25%

Standard Deviation=sqrt(25%*(75000/100000-1-31.25%)^2+50%*(125000/100000-1-31.25%)^2+25%*(200000/100000-1-31.25%)^2)=44.63%

Sharpe Ratio=(31.25%-3%)/44.63%=0.632982299

2.
=(25%*75000+50%*125000+25%*200000)/(1+8%+3%)
=$118,243.24